# Weinstein exactness of nearby Lagrangians and the Lagrangian $C^{0}$ flux conjecture

### v2

@@ 3.2 From rationality to $H$-rationality
Finally, note that, as $v\to 0$, $L_{v}\to K$ in the Hausdorff metric. Therefore, however small we take a neighbourhood of the zero-section of $T^{*}K$, there is a weakly exact Lagrangian in that neighbourhood which is not exact. Therefore, unlike in Theorem 4, there is not an equivalent to Corollary 27 with $\mathcal{L}{\mathrm{Ham}}(L)$ replaced by the space of all $\tau$-rational Lagrangians if $[\pi_{1}(L),\pi_{1}(L)]$ is not finite. In particular, most applications in the introduction do not have equivalents in spaces of $\tau$-rational Lagrangians. That being said, it is entirely possible that Corollary 27 holds in $\mathcal{L}{\mathrm{Ham}}(L)$ without the hypothesis on $[\pi_{1}(L),\pi_{1}(L)]$ — it is however beyond the scope of the techniques presented in this paper. In other words, with the present techniques, it is unclear whether Conjecture B is true for rational Lagrangians in general, even if we restrict to those diffeomorphism types covered by Theorem 3.
4 Proof of Theorem 3
We now turn our attention to Theorem 3. To obtain a proof, we introduce some capacities inspired by work of Cieliebak and Mohnke [CM18] (Section 4.1 below), and we explain how their finiteness implies the theorem (Section 4.2). We conclude by an explicit proof of Theorem 3 in the case of Klein bottles which gives a better estimate in that case (Section 4.3). Note that the methods developed here will also be central to the proof of Theorem 6.
@@ 4.1 Some capacities à la Cieliebak–Mohnke
Let $Q$ be a closed connected $n$-manifold. For any $2n$-dimensional symplectic manifold $(X,\omega)$, we define two classes of Lagrangians:
@@
$\mathscr{L}_{Q}^{0}(X)$ :=$:=\{L=f(Q)\in\mathscr{L}_{Q}(X)\ |\ H_{1}(f)\otimes\mathds{R}=0\},$
where $H_{1}(f)\otimes\mathds{R}$ is the map induced by $f$ on first homology with real coefficients. In turn, this defines two capacities:
$c_{Q}(X)$ $:=\sup\{A^{H}_{\mathrm{min}}(L,X)\ |\ L\in\mathscr{L}_{Q}(X)\}\in[0,+\infty]\quad\mbox{and}$
@@
We take the convention that $c_{Q}(X)=0$ (respectively $c_{Q}^{0}(X)=0$) if $\mathscr{L}_{Q}(X)=\varnothing$ (respectively $\mathscr{L}_{Q}^{0}(X)=\varnothing$). Obviously, we have that $c_{Q}^{0}\leq c_{Q}$. Finally, we set
$c_{\mathrm{all}}(X)$ $:=\sup c_{Q}(X)\quad\mbox{and}$
@@
However, we could develop an entirely analogous theory using homotopy. With it, we would get a version of Theorem 3 for rational Lagrangians, i.e. a neighbourhood of weak exactness. However, this is in general not enough to get a neighbourhood of exactness as in Theorem 4 — see Section 3.2 for a discussion as to when that is the case.
@@
Let $c$ denote $c_{Q}$, $c^{0}_{Q}$, $c_{\mathrm{all}}$, or $c_{\mathrm{all}}^{0}$. We have the two following properties.
@@
$A^{H}_{\mathrm{min}}(L,X)$ $=\inf_{\begin{subarray}{c}u\in H_{2}(X,L;\mathds{Z})\\ \omega(u)>0\end{subarray}}\omega(u)$
$=\inf_{\begin{subarray}{c}a\in\partial H_{2}(X,L;\mathds{Z})\\ \lambda(a)>0\end{subarray}}\lambda(a)$
@@
$=B\cdot\inf_{\begin{subarray}{c}a\in H_{1}(L;\mathds{Z})\\ \lambda(a)>0\end{subarray}}\lambda(a)$
@@
$\leq B\cdot\inf_{\begin{subarray}{c}a^{\prime}\in\partial H_{2}(X^{\prime},\iota(L);\mathds{Z})\\ \lambda^{\prime}(a^{\prime})>0\end{subarray}}\lambda^{\prime}(a^{\prime})$
$=B\cdot A^{H}_{\mathrm{min}}(\iota(L),X^{\prime}).$
@@
We note that $c_{Q}(X)$ and $c^{0}_{Q}(X)$ are equal whenever $H_{1}(X;\mathds{R})=0$. However, there is another important case where they also coincide.
@@
If $\dim H_{1}(Q;\mathds{R})=1$ and $X$ is exact, we have that $c_{Q}(X)=c^{0}_{Q}(X)$.
Proof.
We fix $Q$ and $X$ as above. We can assume that there is a Lagrangian embedding $f:Q\hookrightarrow X$; otherwise both capacities are equal to 0. Since $\dim H_{1}(Q;\mathds{R})=1$, $H_{1}(f)\otimes\mathds{R}$ is either 0 or injective. Suppose that it is injective. By the long exact sequence in homology, we then get that the boundary map $\partial:H_{2}(X,L;\mathds{R})\to H_{1}(L;\mathds{R})$ is zero, where $L=f(Q)$. Since $\omega(H_{2}(X,L;\mathds{R}))=\lambda(\partial(H_{2}(M,L;\mathds{R})))$ whenever $\omega=d\lambda$, we then conclude that $L$ is $H$-exact. In particular, $L\notin\mathscr{L}_{Q}(X)$. Therefore, we have that $\mathscr{L}_{Q}(X)=\mathscr{L}^{0}_{Q}(X)$, which implies the result.
@@
Suppose that $Q^{\prime}$ admits a $H$-exact Lagrangian embedding in $X^{\prime}$. Then, $c_{Q}(X)\leq c_{Q\times Q^{\prime}}(X\times X^{\prime})$. In particular, $c_{\mathrm{all}}(X)\leq c_{\mathrm{all}}(X\times X^{\prime})$ as soon as $X^{\prime}$ admits a $H$-exact Lagrangian.
@@ 4.2 Finiteness of the capacities
Having enunciated the main properties of our capacities, we now explain how one can get Theorem 3 from their finiteness. To do this, we first formulate a meta result.
@@
Let $L$ be a Lagrangian in a symplectic manifold $M$. Suppose that $c_{Q}(D^{*}_{R}L)$ is finite for some $R>0$. For every $\tau\geq 0$, there exists a Weinstein neighbourhood $\mathcal{W}(L)$ of $L$ in $M$, such that all $L^{\prime}\in\mathcal{L}(Q,\tau)$ included in $\mathcal{W}(L)$ is $H$-exact in $\mathcal{W}(L)$.
If $c_{\mathrm{all}}(D^{*}_{R}L)$ is finite for some $R>0$, then the same holds for all $L^{\prime}\in\mathcal{L}(\tau)$.
We recall that $\mathcal{L}(\tau)$ denotes the space of $\tau$-$H$-rational Lagrangians of $M$; $\mathcal{L}(Q,\tau)$ is the subspace of those Lagrangians which have the diffeomorphism type of $Q$.
@@
We only prove the case $c_{Q}(D^{*}_{R}L)<\infty$, as the case $c_{\mathrm{all}}(D^{*}_{R}L)<\infty$ is completely analogous. To do so, we make the following two observations.
  1. (1)
    If $c_{Q}(D^{*}_{R}L)<\infty$, then $\lim_{r\to 0}c_{Q}(D^{*}_{r}L)=0$.
  2. (2)
    If $L^{\prime}\in\mathcal{L}(Q,\tau)$ and $L^{\prime}\subseteq\mathcal{W}_{r}(L)$, then $c_{Q}(D^{*}_{r}L)\geq\tau$ whenever we have that $\omega(H_{2}(\mathcal{W}_{r}(L),L^{\prime}))\neq 0$.
Obviously, the theorem follows directly from these two observations.
The first observation follows directly from Property (i) of Lemma 29. Indeed, we have that
$\lim_{r\to 0}c_{Q}(D^{*}_{r}L)=\lim_{r\to 0}\frac{r}{R}c_{Q}(D^{*}_{R}L)=0.$
Here, we have made use of the fact that $(D^{*}_{r}L,\omega_{0})$ is symplectomorphic to $(D^{*}_{r/a}L,a\omega_{0})$ via the map $(q,p)\mapsto(q,ap)$. Note that Property (ii) of Lemma 29 implies that our capacity is invariant under symplectomorphisms.
For the second observation, take $L^{\prime}\in\mathcal{L}(Q,\tau)$ such that $L^{\prime}\subseteq\mathcal{W}_{r}(L)$, and suppose there is some class $u\in H_{2}(\mathcal{W}_{r}(L),L^{\prime})$ such that $\omega(u)\neq 0$. Without loss of generality, we may suppose that $\omega(u)>0$. By definition of $H$-rationality, there is some $k\in\mathds{Z}$ such that $\omega(A)=k\tau$. Since $\omega(u)>0$, $k\geq 1$. Therefore, $\omega(u)\geq\tau$. Taking the infimum over all possible $u$’s, we get
$\tau\leq A^{H}_{\min}(L^{\prime},\mathcal{W}_{r}(L))\leq c_{Q}(\mathcal{W}_{r}(L))=c_{Q}(D^{*}_{r}L).$
Again, we have made use of the fact that $c_{Q}$ is invariant under symplectomorphisms.
@@
Case $\dim H_{1}(Q;\mathds{R})=1$.
We now turn our attention to the capacity $c^{0}_{Q}$. Recall that from Lemma 31, $c^{0}_{Q}(X)=c_{Q}(X)$ whenever $\dim H_{1}(Q;\mathds{R})=1$ and $X$ is exact. However, studying directly $c^{0}_{Q}$ allows us to show the following general result.

### v3

@@ 3.3 Why $H$-rationality in general?
Finally, note that, as $v\to 0$, $L_{v}\to K$ in the Hausdorff topology. Therefore, however small we take a neighbourhood of the zero-section of $T^{*}K$, there is a weakly exact Lagrangian in that neighbourhood which is not exact. Therefore, there is no homotopical equivalent of Theorem 5 with $\mathcal{L}{\mathrm{Ham}}(L)$ replaced by the space of all $\tau$-rational Lagrangians if $[\pi_{1}(L),\pi_{1}(L)]$ is not finite. In particular, many applications in the introduction do not have equivalents in spaces of $\tau$-rational Lagrangians.
4 Proofs of Theorems 2 and 4
We now turn our attention to Theorem 2. To obtain a proof, we first introduce some capacities inspired by work of Cieliebak and Mohnke [CM18] (Section 4.1), and we explain how their finiteness implies the first half of the theorem (Section 4.2). We show that this straightforwardly yields Theorem 4 (Section 4.3). We then prove that Conjecture C is closed under covering, which concludes the proof of Theorem 2 (Section 4.4). Finally, we state and prove its homotopical variant (Section 4.5). Note that the methods developed here will also be central to the proof of Theorem 7.
@@ 4.1 Some capacities à la Cieliebak–Mohnke
Let $Q$ be a closed connected $n$-manifold. For any $2n$-dimensional symplectic manifold $(X,\omega)$, we define three classes of Lagrangians:
@@
${\mathscr{L}_{Q}^{{}^{\prime}}(X)}$ :=$:=\{L=f(Q)\in\mathscr{L}_{Q}(X)\ |\ \operatorname{\mathrm{im}}(H_{1}(f)\otimes\mathds{R})\neq H_{1}(X;\mathds{R})\mbox{ or }H_{1}(X;\mathds{R})=0\}$
$\mathscr{L}_{Q}^{0}(X)$ :=$:=\{L=f(Q)\in\mathscr{L}_{Q}(X)\ |\ H_{1}(f)\otimes\mathds{R}=0\}$
where $H_{1}(f)\otimes\mathds{R}$ is the map induced by $f$ on first homology with real coefficients.
Lemma 27
We always have the inclusions $\mathscr{L}^{0}_{Q}(X)\subset\mathscr{L}^{{}^{\prime}}_{Q}(X)\subset\mathscr{L}_{Q}(X)$.
Moreover, we have the following particular cases.
  • If $H_{1}(X;\mathds{R})=0$, we have $\mathscr{L}^{0}_{Q}(X)=\mathscr{L}^{{}^{\prime}}_{Q}(X)=\mathscr{L}_{Q}(X)$.
  • If $\dim H_{1}(Q;\mathds{R})=1$ and $X$ is exact, we have $\mathscr{L}^{0}_{Q}(X)=\mathscr{L}^{{}^{\prime}}_{Q}(X)=\mathscr{L}_{Q}(X)$.
  • If $\dim H_{1}(Q;\mathds{R})\leq\dim H_{1}(X;\mathds{R})$ and $X$ is exact, we have $\mathscr{L}^{{}^{\prime}}_{Q}(X)=\mathscr{L}_{Q}(X)$.
Proof.
The general statement and the first point are obvious.
Now, we fix $Q$ and $X$ as in the second point. We can assume that there is a Lagrangian embedding $f:Q\hookrightarrow X$; otherwise all sets are empty. Since $\dim H_{1}(Q;\mathds{R})=1$, $H_{1}(f)\otimes\mathds{R}$ is either 0 or injective. Suppose that it is injective. By the long exact sequence in homology, we then get that the boundary map $\partial:H_{2}(X,L;\mathds{R})\to H_{1}(L;\mathds{R})$ is zero, where $L=f(Q)$. Since $\omega(H_{2}(X,L;\mathds{R}))=\lambda(\partial(H_{2}(M,L;\mathds{R})))$ whenever $\omega=d\lambda$, we then conclude that $L$ is $H$-exact. In particular, $L\notin\mathscr{L}_{Q}(X)$. Therefore, we have that $\mathscr{L}_{Q}(X)=\mathscr{L}^{0}_{Q}(X)$.
Finally, for the third point, $\omega(H_{2}(X,L;\mathds{Z}))\neq 0$ implies by the long exact sequence of a pair that $H_{1}(f)\otimes\mathds{R}$ has non-trivial kernel and therefore $\dim\operatorname{\mathrm{im}}(H_{1}(f)\otimes\mathds{R})\leq\dim H_{1}(X;\mathds{R})-1$.
In turn, this defines three capacities:
$c_{Q}(X)$ $:=\sup\{A^{H}_{\mathrm{min}}(L,X)\ |\ L\in\mathscr{L}_{Q}(X)\}\in[0,+\infty];$
$c_{Q}^{{}^{\prime}}(X)$ $:=\sup\{A^{H}_{\mathrm{min}}(L,X)\ |\ L\in\mathscr{L}_{Q}^{{}^{\prime}}(X)\}\in[0,+\infty];$
@@
We take the convention that $c_{Q}(X)=0$ (respectively $c_{Q}^{0}(X)=0$, $c_{Q}^{{}^{\prime}}(X)=0$) if $\mathscr{L}_{Q}(X)=\varnothing$ (respectively $\mathscr{L}_{Q}^{0}(X)=\varnothing$, $\mathscr{L}_{Q}^{{}^{\prime}}(X)=\varnothing$). Obviously, we have that $c_{Q}^{0}\leq c^{\prime}_{Q}\leq c_{Q}$. Finally, we set
$c_{\mathrm{all}}(X)$ $:=\sup c_{Q}(X),$
${c_{\mathrm{all}}^{{}^{\prime}}(X)}$ $:=\sup c_{Q}^{{}^{\prime}}(X)\quad\mbox{and},$
@@
However, we could develop an entirely analogous theory using homotopy and get some homotopical version of Theorem 4 — see Section 3.2 for a discussion.
@@
Let $c$ denote $c_{Q}$, $c^{\prime}_{Q}$, $c^{0}_{Q}$, $c_{\mathrm{all}}$, $c^{\prime}_{\mathrm{all}}$ or $c_{\mathrm{all}}^{0}$. We have the two following properties.
@@
$A^{H}_{\mathrm{min}}(L,X)$ $=\inf_{\begin{subarray}{c}a\in\partial H_{2}(X,L;\mathds{Z})\\ \lambda(a)>0\end{subarray}}\lambda(a)$
@@
$\leq B\cdot A^{H}_{\mathrm{min}}(\iota(L),X^{\prime}).$
@@
From Lemma 27 above, we directly get.
@@
  • If $H_{1}(X;\mathds{R})=0$, we have $c_{Q}(X)=c^{\prime}_{Q}(X)=c^{0}_{Q}(X)$.
  • If $\dim H_{1}(Q;\mathds{R})=1$ and $X$ is exact, we have $c_{Q}(X)=c^{\prime}_{Q}(X)=c^{0}_{Q}(X)$.
  • If $\dim H_{1}(Q;\mathds{R})\leq\dim H_{1}(X;\mathds{R})$ and $X$ is exact, we have $c^{\prime}_{Q}(X)=c_{Q}(X)$.
@@
Suppose that $Q^{\prime}$ admits a $H$-exact Lagrangian embedding in $X^{\prime}$. Then, $c_{Q}(X)\leq c_{Q\times Q^{\prime}}(X\times X^{\prime})$. In particular, $c_{\mathrm{all}}(X)\leq c_{\mathrm{all}}(X\times X^{\prime})$ as soon as $X^{\prime}$ admits a $H$-exact Lagrangian. The same holds for the corresponding $c^{\prime}$ capacities
@@ 4.2 Finiteness of the capacities
Having enunciated the main properties of our capacities, we now explain how one can get the first part of Theorem 2 from their finiteness.
@@
Let $L$ be a closed connected manifold. There exists $R>0$ such that $c^{{}^{\prime}}_{\mathrm{all}}(D^{*}_{R}L)$ is finite if and only if Conjecture C holds.
@@
Suppose that $c^{{}^{\prime}}_{\mathrm{all}}(D^{*}_{R}L)$ is finite, and set $C:=c^{\prime}_{\mathrm{all}}(D^{*}_{R}L)/R$. It follows from Property (i) of Lemma 28 that
$c^{\prime}_{\mathrm{all}}(D^{*}_{r}L)=\frac{r}{R}c^{\prime}_{\mathrm{all}}(D^{*}_{R}L)=Cr.$
Here, we have made use of the fact that $(D^{*}_{r}L,\omega_{0})$ is symplectomorphic to $(D^{*}_{r/a}L,a\omega_{0})$ via the map $(q,p)\mapsto(q,ap)$. Note that Property (ii) of Lemma 28 implies that our capacity is invariant under symplectomorphisms. Let $K$ be a $H$-rational Lagrangian in $D^{*}_{r}L$, with $H$-rationality constant $\tau_{K}$, such that $(\pi_{K})_{*}$ is not surjective on $H_{1}(\,\cdot\,;\mathds{R})$, i.e. $K$ is in $\mathscr{L}_{K}^{{}^{\prime}}(D^{*}_{r}L)$. Then
$\tau_{K}=A^{H}_{\mathrm{min}}(K,D^{*}_{r}L)\leq c^{\prime}_{\mathrm{all}}(D^{*}_{L})=Cr,$
which is what we wanted to show.
On the other hand, if Conjecture C holds, then $A^{H}_{\mathrm{min}}(K,D^{*}_{R}L)\leq CR$ for all $H$-rational Lagrangians in $D^{*}_{R}L$ such that $(\pi_{K})_{*}$ is not surjective on $H_{1}(\,\cdot\,;\mathds{R})$. Since $A^{H}_{\mathrm{min}}(K,D^{*}_{R}L)=0$ when $K$ is $H$-irrational, we directly get a bound on the capacity. ∎
@@
Let $L_{1}$, …, $L_{k}$ be closed manifolds with $\dim H_{1}(L_{i};\mathds{R})=1$ such that $L_{i}$ embeds as a Lagrangian in a Liouville domain $W_{i}$ such that $SH(W_{i})=0$. Let $L_{0}$ be a closed manifold with $H_{1}(L_{0};\mathds{R})=0$. We have that
$c^{{}^{\prime}}_{\mathrm{all}}\left(D^{*}_{r_{0}}L_{0}\times\dots\times D^{*}_{r_{k}}L_{k}\right)<\infty$
for small enough $r_{i}$’s. Moreover, the same is true for $c_{Q}$ for a closed connected manifold $Q$ with $\dim H_{1}(Q;\mathds{R})\leq k.$
Proof of Theorem 35.
We start with the case $L_{0}$ is a point and remove it from the notation for now. Let $L$ be a Lagrangian in the product of codisk bundles $X$ diffeomorphic to $Q$ and such that $\omega(H_{2}(X,L))\neq 0$. Further suppose that $f_{*}:H_{1}(Q;\mathds{R})\rightarrow H_{1}(X;\mathds{R})$ is not surjective. Denote by
$p_{i}:L_{1}\times\dots\times L_{k}\to L_{1}\times\dots\times\widehat{L_{i}}\times\dots\times L_{k}$
the projection onto the product of all $L_{j}$’s except $L_{i}$ and by $\Pi_{i}$ its lift between products of cotangent bundles.
Since $f_{*}$ is not surjective, its image $I$ is a vector subspace of $H_{1}(X;\mathds{R})$ of dimension less than or equal to $k-1$. Therefore, there exists at least one $i\in\{1,\dots,k\}$ such that, under the identification $H_{1}(X;\mathds{R})\cong\mathds{R}^{k}$, the projection $\pi_{i}:\mathds{R}^{k}\rightarrow R_{i}=\{x\in\mathds{R}^{k}\,|\,x_{i}=0\}$ is injective on $I$. Note that $(\Pi_{i})_{*}$ is naturally identified with $\pi_{i}$. Therefore, for that $i$, we have that $\ker f_{*}=\ker(\Pi_{i}\circ f)_{*}$.
Take $r_{i}$ small enough so that $D^{*}_{r_{i}}L_{i}$ symplectically embeds in $W_{i}$ as in the statement of the theorem. Denote by
$\Psi_{i}:X\hookrightarrow X^{\prime}_{i}:=D^{*}_{r_{1}}L_{1}\times\dots\times W_{i}\times\dots\times D^{*}_{r_{k}}L_{k}$
the symplectic embedding it induces and by $\omega_{i}^{\prime}$ the symplectic form on $X_{i}^{\prime}$. Note that $\ker(\Pi_{i}\circ f)_{*}=\ker(\Psi_{i}\circ f)_{*}$.
Suppose $u^{\prime}\in H_{2}(X_{i}^{\prime},L;\mathds{Z})$ such that $\omega^{\prime}_{i}(u^{\prime})\neq 0$. This implies that $\partial u^{\prime}$ is in $\ker(\Psi_{i}\circ f)_{*}=\ker f_{*}$ in $H_{1}(Q;\mathds{R}$). In particular, there is some $N\in\mathds{Z}$ and some $u\in H_{2}(X,L;\mathds{Z})$ such that $\partial u=N\partial u^{\prime}$. Furthermore, some diagram chasing gives that $N$ must divide the order $B_{i}$ of the torsion part of $H_{i}(D^{*}_{r_{i}}L_{i})$. Moreover, since $\Psi_{i}$ is an exact symplectic embedding, we have $\omega(u)=k\omega^{\prime}(u^{\prime})$. Hence
$A^{H}_{\min}(L,X)\leq B_{i}A^{H}_{\min}(L,X^{\prime}_{i})\leq B_{i}c_{\mathrm{all}}(X^{\prime}_{i})$
Taking the supremum over all possible $L$, we thus get
$c^{{}^{\prime}}_{Q}(X)\leq\max_{i}B_{i}c_{\mathrm{all}}(X^{\prime}_{i}),$
and finiteness follows from Zhou’s theorem.
When $L_{0}$ is not a point, one can check that the entire proof above follows in the same way. ∎
4.3 Proof of Theorem 4
For completeness, we now give a quick proof of how Theorem 4 follows from Conjecture C when it holds.
Proof of Theorem 4.
For all sufficiently small $r>0$, there is a Weinstein neighbourhood $W_{r}(L)\cong D_{r}^{*}L$ of $L$ in $M$. Let $L^{\prime}\in\mathcal{L}(L,\tau)$, and suppose that the map $(\pi|_{L^{\prime}})_{*}:H_{1}(L^{\prime};\mathds{R})\rightarrow H_{1}(L;\mathds{R})$, induced by the restriction of the projection $\pi:D^{*}L\rightarrow L$ to $L^{\prime}$ is not surjective. Then, there is some positive integer $k$ such that $L^{\prime}$ has rationality constant $k\tau$ in $\mathcal{W}_{r}(L)$. If Conjecture C holds, then we have that $k\tau\leq Cr$ for some $C=C(L)>0$. We thus get a contradiction if $r<\tau/C$.
For such $r$, $(\pi|_{L^{\prime}})_{*}:H_{1}(L^{\prime};\mathds{R})\rightarrow H_{1}(L;\mathds{R})$ must thus be surjective whenever $L^{\prime}\subseteq\mathcal{W}_{r}(L)$. Since $H_{1}(L^{\prime};\mathds{R})\cong H_{1}(L;\mathds{R})$, this means that it is an isomorphism, and the result follows from the central lemma (Lemma 24). ∎
We remark that using the methods above and the $c^{0}$ variant of the capacities, we can show the following interesting partial result.
Theorem 36